OpenStudy (anonymous):

multiply and simplify (√(x)-2√(2))(√(x)+2√(2)) (√(x)+√(2))^2

OpenStudy (tkhunny):

Top one - difference of squares. Do these with your eyes closed. \(\sqrt{x}^2 = x\) \((2\sqrt{2})^{2} = 8\) You're almost done.

OpenStudy (anonymous):

this is all for the first equation @tkhunny?

OpenStudy (tkhunny):

That's why I said "top one". (a+b)(a-b) = a^2 - b^2 The bottom one is the square of a binomial. This should give you a perfect square trinomial. You should be familiar with this structure, too. (a+b)^2 = a^2 + 2ab + b^2